THE PFI-ZEKE SPECTRUM OF HfF^{+}, IN SUPPORT OF FUNDAMENTAL PHYSICS

\underline{BEAU J. BARKER}, IVAN O. ANTONOV, VLADIMIR E. BONDYBEY, and MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The HfF^{+} cation has been identified as a molecule with favorable properties for investigation of the dipole moment of the electron. The ion is predicted to have a^{1}Σ^{+} ground state, but the state of greatest interest is the low-lying^{3}Δ_{1} state, which correlates with Hf^{+}(6s5d)F^{−}. A high internal electric field may be generated when the Ω=1 state is polarized by a modest external field. In the present work, spectroscopic data for the ground and low-lying states HfF^{+} have been obtained using the technique of pulse field ionization - zero electron kinetic energy (PFI-ZEKE) spectroscopy. Sequential two-photon excitation was used, with the first photon set to excite a transition near 28593 cm^{−1}. This previously unreported band was used as it is at slightly less than half of the ionization energy (IE), and therefore not subject to one-color, two-photon ionization. PFI-ZEKE spectra were recorded for the levels X^{1}Σ^{+} (v=0-6), ^{3}Δ_{1} (v=0-3), ^{3}Δ_{2} (v=0-3), and ^{3}Δ_{3} (v=0,1). Rotational resolution was achieved using single rotational line excitation of the intermediate state. The IE for HfF was found to be 59477 cm^{−1}. Term energies and molecular constants for the ground and low-lying states of HfF^{+} will be reported.